1)Assume that the total marks were at an average of u=18 per assesment,and that the distribution of total marks are normally distributed with o=10

a)What is the proportion that a student would have a mark more than 24 mark if randomly selected ?

b)What proportion of students would have marks between 10 and 24



2)A normal distribution has u=80 o=10. What is the probability of random selecting the following scores?

a) x > 75

b) x < 85

c) between the mean and score of 90

d) between the mean and score of 50

e) 75 < x <85



1
Expert's answer
2021-06-02T13:04:07-0400

1.)

Given,

"\mu=18\\\sigma=10"


a) Probability that a student have mark more than 24

"P(X>24)"


"z=\dfrac{X-\mu}{\sigma}=\dfrac{24-18}{10}=0.6"


Using Standard normal distribution table:

So, "P(X>24)=P(z>0.6)=0.2743"


b)Probability that a student have marks between 10 and 24


"P(10<X<24)"


"z_1=\dfrac{X-\mu}{\sigma}=\dfrac{10-18}{10}=-0.8\\\ \\z_2=\dfrac{X-\mu}{\sigma}=\dfrac{24-18}{10}=0.6"


Using Standard normal distribution table:




Hence, "P(10<X<24)=P(-0.8<z<0.6)=0.5139"



2.) Given,

"\mu=80\\\sigma=10"


(a) "P(X>75)"


"z=\dfrac{X-\mu}{\sigma}=\dfrac{75-80}{10}=-0.5"


Using Standard normal distribution table:

So, "P(X>75)=P(z>-0.5)=0.6915"




(b) "P(X<85)"


"z=\dfrac{X-\mu}{\sigma}=\dfrac{85-80}{10}=0.5"


Using Standard normal distribution table:

So, "P(X<85)=P(z<0.5)=0.6915"



(c) between the mean and score of 90


"z_1=\dfrac{X-\mu}{\sigma}=\dfrac{80-80}{10}=0"


"z_2=\dfrac{X-\mu}{\sigma}=\dfrac{90-80}{10}=1"


Using Standard normal distribution table:


"P(0<z<1)=0.3413"



(d)  between the mean and score of 50



"z_1=\dfrac{X-\mu}{\sigma}=\dfrac{80-80}{10}=0"


"z_2=\dfrac{X-\mu}{\sigma}=\dfrac{50-80}{10}=-3"


Using Standard normal distribution table:




"P(-3<z<0)=0.4987"



(e) 75 < x <85



"z_1=\dfrac{X-\mu}{\sigma}=\dfrac{75-80}{10}=-0.5"


"z_2=\dfrac{X-\mu}{\sigma}=\dfrac{85-80}{10}=0.5"


Using Standard normal distribution table:




"P(75<X<85)=P(-0.5<z<0.5)=0.3829"